# Neural Network Quantum States for the Interacting Hofstadter Model with Higher Local Occupations and Long-Range Interactions

### v1

@@ I Introduction
Here we show that neural network quantum states (NQS) along with variational Monte Carlo methods provide a promising framework to simulate long-range interacting many-body systems. Such states represent a many-body wave function based on an artificial neural network, as first proposed in a seminal paper by Carleo and Troyer in 2017 [41]. Since then, scientists have worked out many strengths and weaknesses of this ansatz [42, 43, 44, 45, 46]: In recent publications, NQS have been successfully implemented for bosonic systems [47, 48, 49, 50, 51]. In addition, it has been shown to be possible to capture topological phases in general with NQS [52, 53, 54, 55]. A promising ansatz for efficient and accurate network architectures is given by using autoregressive networks, such as recurrent neural networks (RNNs), as they allow to sample directly from the wave function [56, 42, 57, 58, 59].
@@ III.1 From continuum to discrete lattice
First, we consider a square lattice with lattice constant $a$ and vary the magnetic length by tuning the flux per plaquette. We furthermore add on-site Hubbard repulsion, resulting in interacting many-body states similar to previously studied fractional quantum Hall states on the lattice [10, 11, 12, 13, 14, 7, 8, 15, 16, 17, 18, 19, 20, 21, 22, 23] as well as the non-Abelian Pfaffian state [25, 26, 27].
@@ III.4.1 Rydberg-dressing induced interactions
III.4.2 Proposed phase diagram on a lattice
Figure 6: NQS phase diagram for the Hofstadter model with long-range Rydberg-dressed interactions. Marked in blue is a Wigner crystal phase, green is the bubble crystal phase. At $a_{B}\rightarrow 0$ we expect physics from on-site Hubbard-type interactions. To obtain this phase diagram we performed calculations at parameters corresponding to the symbols. The different symbols indicate which regime/phase the NQS ground state belongs to, while the background color is used as a guide to the eye. We used various numbers of particles with $N\in[4,21]$
In this section we discuss the NQS-phase diagram for the simulation of Rydberg-dressed atoms on a lattice with a synthetic magnetic field. The Hamiltonian of the system reads:
@@ III.4.2 Proposed phase diagram on a lattice
The phase diagram that we find for the considered system contains three different regimes, see Fig. 6. The first is the Hofstadter-Bose-Hubbard (HBH) regime, which we obtain for $a_{\mathrm{B}}\rightarrow 0$. We know from the benchmark results in subsection III.3 that this regime can be represented by the RNN-NQS. Further, we find a Wigner crystal phase for intermediate $a_{B}$ and $\nu$ (I), a bubble crystal for large $a_{B}$ (II). Note that there is evidence for a liquid of clustered particles (large $a_{\mathrm{B}}$ and $\nu$) in the continuum  [40, 13, 79, 89], that we do not find on the lattice (III). Additionally, we find first indications of a stripe phase at small filling $\nu$ with a strong magnetic field of $\alpha=0.5$ (see Appendix D.2). To allow a high occupation number for clustered states, we are using a local Hilbert space cut-off of $N_{\mathrm{max}}=3$ (see Tab. 2). In the following, we discuss the phases (I-III) in more detail.
@@
with $i=(x_{i},y_{i})$ and $j=(x_{j},y_{j})$. In order to obtain $g^{(2)}(r)$ we average over all possible combinations of $i$ and $j$ with distance $r$. The maximum error is calculated by error propagation of the standard error of the mean: $\Delta\langle\hat{n}_{i}\hat{n}_{j}\rangle$, $\Delta\langle\hat{n}_{i}\rangle$ and $\Delta\langle\hat{n}_{j}\rangle$.
The error for $g^{2}(r=0)$ is often quite large, because $\langle\hat{n}_{i}\hat{n}_{i}\rangle-\langle\hat{n}_{i}\rangle\lesssim\Delta\langle\hat{n}_{i}\hat{n}_{j}\rangle+\Delta\langle\hat{n}_{i}\rangle$ with $\langle\hat{n}_{i}\hat{n}_{j}\rangle\gg\Delta\langle\hat{n}_{i}\hat{n}_{j}\rangle$ and $\langle\hat{n}_{i}\rangle\gg\Delta\langle\hat{n}_{i}\rangle$.
@@
(i) It is likely that there is indeed no cluster liquid for the considered lattice system, as the predefined lattice of the system in combination with the large blockade radius may make free movement of clusters unfavorable. Further, we observe that the crystal lattices of bubble crystals, which compete with cluster liquids, are incommensurable with the system size. E.g. the density profile in Fig. 8 cannot be used as a building block for larger systems, but would require padding of three empty sites in the x and y direction. Hence, bubble crystals lower their energy in finite-size systems by increasing the effective system size ($\nu_{\mathrm{eff}}<\nu$). This amounts to strong finite-size corrections at the considered system sizes.
@@ Appendix C Additional details on the benchmark
To obtain additional information about our results, we compare the relative error and the normalized variances of the energy. The normalized quantum mechanical variance of the $12\times 12$ benchmark can be seen in Fig. 11. Here, we observe that the variance in the regime (I) ($\alpha=0.1$ and $\alpha\gtrsim 0.35$) is smaller than for the intermediate regime (II). In Fig. 12, selected points from the $6\times 6$ and $12\times 12$ benchmarks are shown. Here, we show the relative error and the normalized statistical variance as a function of the epoch. Note that the statistical variance is evaluated with PyTorch’s internal function that calculates: $\sigma(E)=(\langle\hat{H}^{2}\rangle-|\langle\hat{H}\rangle|^{2})$.
Figure 12: A comparison of exemplary training procedures for a Hofstadter model on a $6\times 6$ ($N=4$ particles) and $12\times 12$ ($N=16$ particles) square lattice with on-site interactions. This figure shows the relative error in the upper and the normalized statistical variance of the energy in the lower picture
Interestingly, the graphs for the $6\times 6$ and for the $12\times 12$ lattice are similar in their structure and convergence speed for 200000 epochs. For $\alpha=0$, both systems converge comparatively fast to the ground state. The other cases take significantly longer to get to the ground state. Note that the variance for $\alpha=0.4$ is lower than for $\alpha=0.25$, although the error is approximately the same. On the $6\times 6$ lattice, the higher variance ($\alpha=0.25$), is caused by a larger gap between the contributing ground and first excited state.
@@ Appendix D Additional Data
D.1 Topologically ordered states
@@
D.2 Stripe phase
Figure 14: Density profile of a typical stripe order with the full counting statistics of particle numbers in each square (panel a)). Panel b) shows the two particle correlator $g^{(2)}(r)$. The parameters are: $N=7$, $\alpha=0.5$, $a_{\mathrm{B}}=1.0$

### v3

@@ I Introduction
Here we show that neural network quantum states (NQS) along with variational Monte Carlo methods provide a promising framework to simulate long-range interacting many-body systems. Such states represent a many-body wave function based on an artificial neural network, as first proposed in a seminal paper by Carleo and Troyer in 2017 [41]. Since then, scientists have worked out many strengths and weaknesses of this ansatz [42, 43, 44, 45, 46]: In recent publications, NQS have been successfully implemented for bosonic systems [47, 48, 49, 50, 51, 52, 53]. In addition, it has been shown to be possible to capture topological phases in general with NQS [54, 55, 56, 57]. A promising ansatz for efficient and accurate network architectures is given by using autoregressive networks, such as recurrent neural networks (RNNs), as they allow to sample directly from the wave function [58, 42, 59, 60, 61].
@@ III.1 From continuum to discrete lattice
First, we consider a square lattice with lattice constant $a$ and vary the magnetic length by tuning the flux per plaquette. We furthermore add on-site Hubbard repulsion, resulting in interacting many-body states similar to previously studied fractional quantum Hall states on the lattice [10, 11, 12, 13, 14, 7, 8, 15, 16, 17, 18, 19, 20, 21, 22, 23, 79] as well as the non-Abelian Pfaffian state [25, 26, 27]. Other exotic states were found in numerical studies of lattice models with Chern numbers $|C|>1$ [80].
@@ III.4.1 Rydberg-dressing induced interactions
III.4.2 Proposed phase diagram for systems on a finite-size lattice
Figure 6: Finite-size phase diagram for the Hofstadter model with long-range Rydberg-dressed interactions obtained with NQS calculations. Marked in blue is a Wigner crystal phase, green is the bubble crystal phase. At $a_{B}\rightarrow 0$ we expect physics from on-site Hubbard-type interactions. To obtain this phase diagram we performed calculations at parameters corresponding to the symbols. The different symbols indicate which regime/phase the NQS ground state belongs to, while the background color is used as a guide to the eye. We used various numbers of particles with $N\in[4,21]$ for the $12\times 12$ system.
In this section we discuss the NQS-phase diagram for the simulation of Rydberg-dressed atoms on a finite-size lattice with a synthetic magnetic field. The Hamiltonian of the system reads:
@@ III.4.2 Proposed phase diagram for systems on a finite-size lattice
The phase diagram that we find for the considered finite-size system contains three different regimes, see Fig. 6. The first is the Hofstadter-Bose-Hubbard (HBH) regime, which we obtain for $a_{\mathrm{B}}\rightarrow 0$. We know from the benchmark results in subsection III.3 that this regime can be represented by the RNN-NQS. Further, we find a Wigner crystal phase for intermediate $a_{B}$ and $\nu$ (I), a bubble crystal for large $a_{B}$ (II). Note that there is evidence for a liquid of clustered particles (large $a_{\mathrm{B}}$ and $\nu$) in the continuum  [40, 13, 82, 92], that we do not find on the lattice (III). Additionally, we find first indications of a stripe phase at small filling $\nu$ with a strong magnetic field of $\alpha=0.5$ (see Appendix D.3). To allow a high occupation number for clustered states, we are using a local Hilbert space cut-off of $N_{\mathrm{max}}=3$ (see Tab. 2). In the following, we discuss the phases (I-III) in more detail.
@@
with $i=(x_{i},y_{i})$ and $j=(x_{j},y_{j})$. In order to obtain $g^{(2)}(r)$ we average over all possible combinations of $i$ and $j$ with distance $r$. The error is estimated using the Jackknife resampling technique.
@@
Note that the boundary conditions stabilize the crystalline phases. Nonetheless, there is evidence of those symmetry broken phases in systems with no boundary effects. This is addressed in the Appendix D.4.
@@
(i) It is likely that there is indeed no cluster liquid for the considered finite-size lattice system, as the predefined lattice of the system in combination with the large blockade radius may make free movement of clusters unfavorable. Further, we observe that the crystal lattices of bubble crystals, which compete with cluster liquids, are incommensurable with the system size. E.g. the density profile in Fig. 8 cannot be used as a building block for larger systems, but would require padding of three empty sites in the x and y direction. Hence, bubble crystals lower their energy in finite-size systems by increasing the effective system size ($\nu_{\mathrm{eff}}<\nu$). This amounts to strong finite-size corrections at the considered system sizes. Note that its not a rare phenomenon that the incommensurability of the filling significantly impacts the physics [93, 94, 95, 96, 97, 98, 99].
@@ IV Conclusion and outlook
Code and data availability
Code and data are available from the corresponding author on request.
@@ Appendix C Additional details on the benchmark
To obtain additional information about our results, we compare the relative error and the normalized variances of the energy. The normalized quantum mechanical variance of the $12\times 12$ benchmark can be seen in Fig. 11. Here, we observe that the variance in the regime (I) ($\alpha=0.1$ and $\alpha\gtrsim 0.35$) is smaller than for the intermediate regime (II). In Fig. 12, selected points from the $6\times 6$ and $12\times 12$ benchmarks are shown. Here, we show the relative error and the normalized statistical variance as a function of the epoch. Note that the statistical variance is evaluated with PyTorch’s internal function that calculates: $\sigma^{2}(E)=(\langle\hat{H}^{2}\rangle-|\langle\hat{H}\rangle|^{2})$.
Figure 12: A comparison of exemplary training procedures for a Hofstadter model on a $6\times 6$ ($N=4$ particles) and $12\times 12$ ($N=16$ particles) square lattice with on-site interactions. This figure shows the relative error in the upper and the normalized statistical variance of the energy in the lower picture.
Interestingly, the graphs for the $6\times 6$ and for the $12\times 12$ lattice are similar in their structure and convergence speed for 200000 epochs. For $\alpha=0$, both systems converge comparatively fast to the ground state. The other cases take significantly longer to get to the ground state. Note that the variance, for both system sizes, is lower for $\alpha=0.4$ than for $\alpha=0.25$, although the error is approximately the same. On the $6\times 6$ lattice, the lower variance of the NQS for $\alpha=0.4$ can be explained by the, on average, smaller energetic differences of the contributing ground and first/second excited state (see Fig. 5). The NQS for $\alpha=0.25$ ($6\times 6$) has a much larger ground state contribution but a higher variance due to the significantly larger energy difference between ground state and first excited state.
@@ Appendix D Additional Data
D.1 Representational power of RNN-NQS
In the benchmark section III.3, we observed that the NQS did not always capture the full ground state (compare to Figure 5). To show evidence that the NQS did not reach its representational limit, we show an exemplary fidelity optimization of the $6\times 6$ system with a magnetic flux of $\alpha=0.28$ in Figure 13. Already after a few thousand epochs, the ground state overlap surpasses the measured overlap of the NQS with the minimization according to the energy gradient ($\sim 89.7\%$ after 200K epochs). We stopped the training after 25K epochs, where we find an ground state overlap of approximately $98\%$. Based on these findings we expect a perfect ground state overlap after further training. Hence, we find clear evidence that the optimization according to the energy gradient is the limiting factor, while the representational limit of the RNN-NQS is not reached for the shown results in section III.3.
Figure 13: Fidelity optimization. We show the infidelity as a function of epochs for the interacting Hofstadter model on a $6\times 6$ lattice. The parameters are: $N=4$, $\alpha=0.28$, $N_{\mathrm{max}}=2$, $U=4$.
D.2 Topologically ordered states
@@
D.3 Stripe phase
Figure 15: Density profile of a typical stripe order with the full counting statistics of particle numbers in each rectangle (panel a)). Panel b) shows the two particle correlator $g^{(2)}(r)$. The parameters are: $N=7$, $\alpha=0.5$, $a_{\mathrm{B}}=1.0$
@@ D.3 Stripe phase
D.4 Symmetry broken phases in systems without stabilizing boundary effects
Figure 16: Density profile and the full counting statistics of particle numbers in each square (panel a)) of a crystalline structure in a system with periodic boundaries. Panel b) shows the two particle correlator $g^{(2)}(r)$. The parameters are: $N=14$, $\alpha=0.25$, $a_{\mathrm{B}}=1.0$
In the subsection III.4.2, we identified two crystalline phases for the considered long-range interacting system. However, these symmetry broken phases often lower their energy by increasing the effective system size, which can make them favorable compared to liquid-like phases.
Nonetheless, there is evidence that the symmetry broken phases persist in systems where the crystalline structure is not stabilize by the boundary conditions. In Figure 16, we show a crystalline phase in a system with periodic boundary conditions. The respective density profile shows 9 crystal sites for 14 particles. Combined with the FCS that indicates $1-2$ particles per crystal lattice site, we can infer that there are on average 5 crystal sites double occupied. The crystal sites have a periodicity of $4$, which is also supported by the two particle correlator that shows a small maximum at $r=4$.
The evidence of crystalline phases in systems with periodic boundary conditions, indicates that we can expect symmetry broken phases in systems with less stabilizing boundary effects.
Note that using PBCs the ground state might be a superposition of crystalline configurations related by translations. However, due to our implementation of the sampling, our simulations favor variational ground states with a high occupation at the first few lattice sites. Nonetheless, there is remaining evidence of a potential superposition, since the full counting statistic is less peaked around a certain particle number.